1. There is a fruit seller who has 30 Kgs of apples to be sold and he wants to fix a price so that all the apples are sold. There are three customers in the market and their individual demand functions are given below:
D1=25-.05P
D2=20-.025P
D3=15-.075P
Where D is the demand and P is the price
1
Expert's answer
2015-10-08T00:00:44-0400
1) If Qs = 30 Kgs of apples to be sold and
there are three customers in the market and their individual demand functions are given below:
D1=25-.05P
D2=20-.025P
D3=15-.075P
Then the market demand equation for the fruit seller is: D = D1 + D2 + D3 = 60 ­ 0.15P
D = S, so:
60 - 0.15P = 30,
0.15P = 30,
P = $200 is the price at which he can sell all the apples.
The individual demands will be:
D1 = 25 - .05*200 = 15 units
D2 = 20 - .025*200 = 15 units
D3 = 15 - .075*200 = 0 units